# Breaking the False Sense of Security in Backdoor Defense through Re-Activation Attack

### v1

@@ 1 Introduction
The pervasive application of Deep Neural Networks (DNNs) across safety-critical domains like facial recognition and autonomous driving [19, 29] has underlined their significance and profound impact in industrial and academic spheres. Despite their transformative potential, DNNs are known to be vulnerable to malicious threats [5, 23], which compromise the integrity and reliability of advanced systems. One of the representative threats is backdoor attacks [15, 27], where an adversary pre-defines a "trigger" and embeds it within limited training data such that the backdoored model will misclassify trigger-containing inputs into specific target categories while appropriately processing benign inputs.
@@ 4.2 Main results
BadNets [15] 93.79 2.01 95.43 1.96 98.42 4.48 97.90 99.17 97.42 1.30 90.17 1.46 96.21
Blended [10] 99.76 99.76 100.00 47.64 99.98 26.83 99.83 98.64 99.63 5.20 93.36 0.20 24.07
Input-Aware [34] 99.30 0.70 99.98 0.92 99.98 0.02 99.77 96.92 21.78 1.26 15.56 0.00 95.28
LF [55] 99.06 99.06 99.93 75.47 99.84 11.99 98.35 93.85 99.04 2.49 96.62 5.43 80.09
SSBA [27] 97.07 97.07 99.27 70.77 99.38 2.89 20.44 98.06 91.29 2.16 95.03 0.54 76.21
Trojan [30] 99.99 2.76 99.76 5.77 99.09 0.54 96.18 96.57 89.58 1.39 83.67 8.93 21.79
WaNet [35] 98.90 98.90 99.72 0.73 99.86 0.88 83.79 98.90 94.67 0.82 89.49 0.26 98.69
Avg 98.26 57.18 99.16 29.04 99.51 6.80 85.18 97.44 84.77 2.09 80.55 2.40 70.33

### v2

@@ 1 Introduction
The pervasive application of Deep Neural Networks (DNNs) across safety-critical domains like facial recognition and autonomous driving [19, 31] has underlined their significance and profound impact in industrial and academic spheres. Despite their transformative potential, DNNs are known to be vulnerable to malicious threats [5, 23, 29, 28], which compromise the integrity and reliability of systems. One of the representative threats is backdoor attacks [15, 27], where an adversary pre-defines a "trigger" and embeds it within limited training data such that the backdoored model will misclassify trigger-containing inputs into specific target categories while appropriately processing benign inputs.
@@ 4.2 Main results
BadNets [15] 93.79 2.01 95.43 1.96 98.42 4.48 97.90 1.63 97.42 1.30 90.17 1.46 96.21
Blended [10] 99.76 99.76 100.00 47.64 99.98 26.83 99.83 12.17 99.63 5.20 93.36 0.20 24.07
Input-Aware [36] 99.30 0.70 99.98 0.92 99.98 0.02 99.77 1.07 21.78 1.26 15.56 0.00 95.28
LF [58] 99.06 99.06 99.93 75.47 99.84 11.99 98.35 6.43 99.04 2.49 96.62 5.43 80.09
SSBA [27] 97.07 97.07 99.27 70.77 99.38 2.89 20.44 4.06 91.29 2.16 95.03 0.54 76.21
Trojan [32] 99.99 2.76 99.76 5.77 99.09 0.54 96.18 4.12 89.58 1.39 83.67 8.93 21.79
WaNet [37] 98.90 98.90 99.72 0.73 99.86 0.88 83.79 0.96 94.67 0.82 89.49 0.26 98.69
Avg 98.26 57.18 99.16 29.04 99.51 6.80 85.18 4.35 84.77 2.09 80.55 2.40 70.33